OpenStudy (anonymous):

why is the sign for the sum and difference formula for cos(u+/-v) = cosucosv-/+sinusinv i.e. the sign is negative instead of positive

OpenStudy (anonymous):

like in calulus when you want to find the derivative of cosx using first principle and the identity is cosucosv-sinusinv

OpenStudy (anonymous):

Are you familiar with Euler's Formula? \[e^{i\theta} = \cos(\theta)+isin(\theta)\] There is a pretty good proof using that fact that im uploading in a sec

OpenStudy (anonymous):

Here ya go :)

OpenStudy (anonymous):

if theres something i put that isnt clear, feel free to ask so i can clarify :)

OpenStudy (anonymous):

no am not familiar with euler's

OpenStudy (anonymous):

i dont understand lol...

OpenStudy (anonymous):

thanks though joe!

OpenStudy (anonymous):

Good luck with that one...the other proof isn't all that great as well.